### v1

# ARTICLE TYPE PAGURI: a user experience study of creative interaction with text-to-music models

Abstract
In recent years, text-to-music models have been the biggest breakthrough in automatic music generation. While they are unquestionably a showcase of technological progress, it is not clear yet how they can be realistically integrated into the artistic practice of musicians and music practitioners. This paper aims to address this question via Prompt Audio Generation User Research Investigation (PAGURI), a user experience study where we leverage recent text-to-music developments to study how musicians and practitioners interact with these systems, evaluating their satisfaction levels. We developed an online tool through which users can generate music samples and/or apply recently proposed personalization techniques, based on fine-tuning, to make the text-to-music model generate sounds closer to their needs and preferences. Using questionnaires, we analyzed how participants interacted with the proposed tool, to understand the effectiveness of text-to-music models in enhancing users’ creativity. Results show that even if the audio samples generated and their quality may not always meet user expectations, the majority of the participants would incorporate the tool in their creative process. Furthermore, they provided insights into potential enhancements for the system and its integration into their music practice.
Keywords:
Up to six keywords (optional).
1. Introduction
Since the 1950s, computer-based music generation has been an interest of both the music and computer science research communities (Hiller Jr and Isaacson,, 1957; Mathews et al.,, 1969). The advent of deep learning has brought immense advancements, substituting all previously existing state-of-the-art techniques (Briot and Pachet,, 2020). In this context, the latest breakthrough has been the recent introduction of Text-To-Music (TTM) models, which enable the generation of raw audio musical signals given as input text prompts describing the desired music.
The introduction of TTM models has lowered the technical competencies needed to use music generative models, effectively posing, for the first time, the need to consider if and how AI is introducible into music creative practice. Although generative models for music are gaining more and more popularity, there is still a lack of comprehensive research on this topic. We advocate that it is important to conduct this type of research closely with potential final users and music practitioners, to develop tools that not only showcase the impressive capabilities of the technology in the music field but also highlight how these tools can be viewed and perceived as musical instruments for creating music.
In this paper, we propose Prompt Audio Generator User Research Investigation (PAGURI). Through PAGURI, we aim to analyze to what extent TTM models are ready to be used as tools for music creation and composition and explore their potential integration into the music creation process based on feedback from users and music professionals. Our main focus is determining the specific phases and purposes for which users would apply these tools in their creation processes, and where they are considered most useful.
We developed an interface through which the user can indicate through a text prompt the audio sample they wish to generate. Additionally, the user can upload up to desired audio samples to personalize the TTM generative model. For the study, we use AudioLDM2 (Liu et al., 2023b, ) as a generative model, and a TTM personalization technique proposed in (Plitsis et al.,, 2023) to let the users fine-tune the model according to the music samples of their choice. We then conducted a user experience study, both online and in presence, where users can interact with PAGURI and we analyzed their experience using questionnaires and open questions. Specifically, we first quantify their background related to music and AI tools, and then we analyze their level of satisfaction after each use of the TTM model, until reaching the desired result. Finally, we let them answer a questionnaire analyzing the whole experience and gathering information related to the perceived usability of the TTM model in music practice. We consider the findings significant contributions to the research community, serving as valuable guidelines for structuring user experience tests involving generative models. Additionally, they represent an initial step in fostering practical discussions regarding user interactions with such models, besides being guidelines for the development and evaluation of TTM models. Supplementary material containing the answers to all questions and feedback is available on the accompanying website111https://paguri-ismir-2024.github.io/PAGURI/.
2. Background
In this section, we provide the reader with the necessary background related to text-based generative music models and techniques used to personalize them.
2.1 Text-to-music models
Language models (Vaswani et al.,, 2017) provide extremely efficient capabilities in generating content with long-term context. As such, they have increasingly become the backbone of recently proposed generative models. Many works have applied them to the task of raw audio music and audio synthesis, employing different types of architectures. The first proposed model was AudioLM (Borsos et al.,, 2023), a multi-stage transformer-based model operating on tokens extracted from audio via SoundStream (Zeghidour et al.,, 2021), and from text via w2v-BERT (Chung et al.,, 2021). This methodology was further developed in MusicLM (Agostinelli et al.,, 2023), where the joint music-text embedding model MuLan (Huang et al.,, 2022) was applied to overcome data scarcity. Similar transformer-based approaches were proposed for general audio synthesis using auto-regressive AudioGen (Kreuk et al.,, 2022), MusicGen (Copet et al.,, 2024), as well as non-auto-regressive models such as MAGNeT (Ziv et al.,, 2024). Several diffusion-based models were also proposed. The first one, DiffSound (Yang et al.,, 2023) considered audio generation by proposing a text-conditioned diffusion model operating on tokenized audio. Later on, several diffusion-based approaches were proposed also for text-to-music generation, such as Make-an-Audio (Huang et al.,, 2023), AudioLDM (Liu et al., 2023a, ), and its evolution AudioLDM2 (Liu et al., 2023b, ).
Question | 1 | 2 | 3 | 4 | 5 |
1) I spend a lot of my free time doing music-related activities | 0 % | 8.3% | 25% | 41.7% | 25% |
2) I enjoy writing about music, for example on blogs and forums | 20.8 % | 54.2% | 16.7% | 8.3% | 0% |
3) If somebody starts singing a song I don’t know, I can usually join in | 12.5 % | 12.5% | 33.3% | 33.3% | 8.3% |
4) I can sing or play music from memory | 0 % | 4.2% | 8.3% | 41.7% | 45.8% |
5) I often read or search the internet for things related to music | 0 % | 0% | 25% | 16.7% | 58.3% |
6) I am able to identify what is special about a given musical piece | 4.2 % | 4.2% | 12.5% | 41.7% | 37.5% |
7) I often read or search the internet for things related to AI tools | 16.7 % | 20.8% | 29.2% | 29.2% | 4.2% |
8) AI can be a tool to support human activities | 0 % | 0% | 25% | 25% | 50% |
2.2 Personalization
Personalization of a model involves adjusting pre-trained generative models to specific examples that the model has not encountered before, as they were not included in the original dataset. In the context of language-based generative models, the task was first introduced for images in (Gal et al.,, 2023), where the authors proposed the Textual-Inversion technique. This technique allows injecting into the model information from data that it did not encounter during training. This is performed by using a few images whose text-embedding representations, or pseudoword (i.e. a token), are learned in the embedding space of the text encoder. Anyway, this technique is limited by the expressive capabilities of the pre-trained generative model. To overcome this limitation, the DreamBooth approach was proposed in (Ruiz et al.,, 2023). In this case, the desired examples are represented using rare token identifiers, and, consequently, the considered generative model (usually diffusion-based) is fine-tuned to learn the desired content.
Later on, personalization techniques were further developed for TTM generative models. This advancement was initially showcased in (Plitsis et al.,, 2023), where Textual Inversion and DreamBooth techniques were integrated with AudioLDM (Liu et al., 2023a, ). The objective was to enhance the model’s ability to learn new music samples and conduct style transfer. Given the potential of incorporating personalization techniques, we decided to use AudioLDM2 in this study.
2.3 Interaction Studies
The interaction of users with text-based generative models has been extensively studied in the image domain (Feng et al.,, 2023; Brade et al.,, 2023), where also datasets of TTI prompts with corresponding preferences given by real users over AI-generated images (Kirstain et al.,, 2024) have been released.
In the audio/music domain, previous research has explored the perception of AI-generated music (Chu et al.,, 2022) and the opportunities and challenges of AI for music creation from a human perception (Newman et al.,, 2023). Several interfaces have been proposed to enable the interaction between users and generative models (Simon et al.,, 2008; Huang et al.,, 2019; Louie et al.,, 2020; Rau et al.,, 2022; Zhang et al.,, 2021; Zhou et al.,, 2020, 2021; Yakura and Goto,, 2023). Among these, the work closest to the one proposed in this paper is IteraTTA (Yakura and Goto,, 2023), an interface for TTM generation enabling iterative exploration of prompts and audio priors. Differently from PAGURI, IteraTTA did not contain the possibility of personalizing the model using audio samples of choice. Moreover, our goals and design strategy are different. In IterATTA the objective was mainly to design an interface for supporting novice users, analyzing user behavior ex-post. In our study, instead, the simple PAGURI interface was designed to study how music practitioners interacted with these models. For this reason, the whole user experience test was conducted live with a more structured interview procedure based on questionnaires.
3. Study design and Method
In this section, we are going to present the study design and the methodology proposed. In section 3.1 we introduce the PAGURI interface, while Section 3.2 presents the procedure followed to conduct the interactive experiment.
3.1 PAGURI Interface
The Prompt Audio Generator User Research Investigation (PAGURI) proposed in this paper presents an interface through which users can generate audio samples using state-of-the-art TTM models. Through the interface, users can generate audio samples giving textual input prompts of their choice to the generative model. They can specify the number of audio tracks to generate and their duration, for each iteration. Additionally, through the same interface, users can upload up to preferred audio files to personalize AudioLDM2. Subsequently, they can obtain a personalized model using the DreamBooth personalization technique as in (Plitsis et al.,, 2023). In this case, the user can specify an instance word used as an identifier to be attached to the new audio files, and an object class which indicates a more general description of the audio, which may indicate the instrument, genre, or musical style of the concepts. Each user can select to either personalize the off-the-shelf AudioLDM2 model or a previously personalized (by the same user) version of it. The fine-tuning duration can be chosen between Fast, Medium and Slow. Each of them applies the DreamBooth personalization procedure for different durations, ranging from a minimum of minutes (Fast) to a maximum of minutes per iteration (Slow). While longer training times enable to obtain better results, we constrained the maximum considered time to retain the interactivity aspect of the study. PAGURI interface has been developed using Python programming language, with the use of Jupyter Notebook and ipywidgets library222https://github.com/jupyter-widgets/ipywidgets. A detailed presentation and images showing PAGURI interface can be found on the supplementary material1.
3.2 Procedure
The procedure was the same for both in-person and online experiments. It can be divided into three substeps:
1.
Preliminary analysis. The participants were introduced to the study and they were required to answer a brief questionnaire containing demographic questions (age, nationality, etc.). To frame the knowledge of participants concerning TTM models, we asked them to answer a second questionnaire regarding their musical knowledge and experiences with AI tools, using questions partially taken from the Goldsmith Musical Sophistication Index (Gold-MSI) (Müllensiefen et al.,, 2014).
2.
Text-To-Music interaction. Participants were asked to interact with the PAGURI interface, generating audio samples and personalizing the TTM model to suit their preferences. After each generation iteration, participants were asked to complete the Model Evaluation Survey1, expressing their satisfaction with the generated audio samples regarding their consistency with respect to the input prompt, audio quality, and alignment with their general expectations.
3.
Final analysis. Upon completion of the experiment, participants were requested to complete a questionnaire regarding their satisfaction with the entire interaction experience with the TTM model via PAGURI. We also asked for open-answer comments and suggestions regarding possible applications and inclusion of TTM models in artistic practice.
The whole procedure lasted approximately one hour. The number of iterations that the participant had with the models during phase 2) was variable and chosen by the users according to their desire. The entire questionnaire forms can be found on the supplementary material1.
4. Results
This section initially focuses on analytics data, followed by the presentation of the main findings from the study.
Question | 1 | 2 | 3 | 4 | 5 |
1) The generated audio(s) is consistent with respect to the input prompt | 18.8 % | 18.8% | 20.7% | 28.9% | 12.8% |
2) The audio(s) generated is consistent with respect to my expectations | 9.9 % | 16.4% | 27.3% | 28.9% | 17.4% |
3) The quality of the audio(s) generated is consistent with my expectations | 14.8 % | 19.1% | 20.7% | 28.6% | 16.8% |
4.1 Demographics Analysis of Participants
A total of people participated to the study, with an average age of years (sd = 5.01). identified as He/Him, while identified as She/Her. To guarantee anonymity, each participant has been assigned an ID. 95% of the participants were of Italian nationality. In terms of current occupation, 62.5% of the participants were master’s degree students, (mainly from the Master of Science in Music and Acoustic Engineering from Politecnico di Milano), 4.2% were bachelor’s degree students, while the remaining 33.3% were workers. Most participants are currently active in various musical endeavors: of them are part of a band or an orchestra, are music producers/mastering engineers or DJs, 2 of them are dance teachers and one is a record label manager. 62.5% of the participants conducted the experiments in person, while 37.5% participated remotely via the Zoom platform.
4.2 Participants Musical knowledge and AI Tool Experience: Demographics and Analysis
The first part of the experiment focused on understanding the relationship of participants with AI tools and music in general. To gather this insight, participants were asked to complete a questionnaire. In Table 1, we present questions where responses were rated on a Likert scale from 1 to 5. For other responses, we provide details in the text. Most of the participants have a strong interest in music, with listening to music daily for more than one hour. engaged in regular, daily practice of an instrument for more than years. At the peak of their interest, more than of the participants practiced for one hour or more daily. Among the participants, of them can play at least one instrument (among which guitar, piano, flute, drums, bass guitar, trombone, ocarina) or are singers.
Regarding the relationship between users and AI tools, is familiar with AI tools and the was able to mention some of them such as ChatGPT (Wu et al.,, 2023), Copilot (Bird et al.,, 2022), Dall-E (Ramesh et al.,, 2021), and Gemini (Team et al.,, 2023). Some TTM tools were also reported, such as Riffusion (Forsgren and Martiros,, 2022), and MusicGen (Agostinelli et al.,, 2023). Focusing on TTM tools, the is aware of their existence, although most of them were unable to explicitly name examples of TTM AI tools. However, while the majority of the participants regularly employ artificial intelligence tools, only the of participants use TTM models.
4.3 Text-To-Music interaction
In total, 304 textual prompts (mean = 12.6 prompts per user, sd = 4.56) were given to the TTM model, for a total of 1148 audio files generated (mean = 47.8 audio files per user). Additionally, there were 39 personalizations made by participants (mean = 1.63 per user), meaning 39 customizations of one of the AudioLDM2 models (either the original AudioLDM2 or a previously personalized model) were done using the DreamBooth technique (Plitsis et al.,, 2023; Ruiz et al.,, 2023). Different audio files were used as input to personalize the model: electric guitar riffs, one-shot bass or synth lead samples, and 8-10 second tracks of musical beats or songs, among others.
When considering textual prompts, it was noted that many users referred to famous people, musical groups, objects of daily use, or cartoon characters, expecting as a response a set of audio tracks somehow connected to them. Most of the time, they did not get the desired response. As an example, P13, in the first iteration, asked the model to generate music related to the prompt daft punk style song, expecting electronic music in the style of the electronic French touch duo Daft Punk. However, the generated sound was rather reminiscent of punk rock. In the same context, in the th consecutive generation of the same user, in response to the request linkin park, audio tracks containing sounds of nature or related to the setting of a natural park were generated instead of music and performances similar to the rock group Linkin Park. Differently, generated audio files related to textual prompts associated with sound effects and soundtracks were highly appreciated by participants.
P2 commented: ”It is necessary to start from easy prompt to understand how to interact with the model, to then move to more complicated and sophisticated prompts”. Most prompts referred to acoustic and electric guitar sound generation, drum beats, piano, and music styles. Also, animal and everyday sounds (such as a cowbell or dog barking) were requested from a high percentage of participants. Some users requested more abstract sounds such as P22: underwater bossa nova classical guitar, and P24: the sound of a drum in space. Others asked for specific sounds, such as one of the prompts of P24: the sound of a 56k modem.
Question | 1 | 2 | 3 | 4 | 5 |
1) I enjoyed the interaction with the text-to-music generation system | 0 % | 0% | 0% | 37.5% | 62.5% |
2)The data generated by the model are consistent with respect to the desired audio file provided by the user* | 4.5 % | 0.0% | 31.8% | 45.5% | 18.2% |
3) The waiting time of the fine-tuning of the model is proportionate with the quality of the generated audio* | 0.0 % | 9.1% | 18.2% | 40.9% | 31.8% |
4) The audio generated by the personalized model is better than the audio generated by the based model* | 4.5 % | 18.2% | 27.3% | 13.6% | 36.4% |
5) There is consistency between input prompt and audio(s) generated | 4.2 % | 4.2% | 33.3% | 58.3% | 0% |
6) The use of text-to-music models can support musicians in musical creation endeavours | 4.2 % | 4.2% | 20.8% | 29.2% | 41.7% |
7) I would use this system again | 4.2 % | 12.5% | 12.5% | 33.3% | 37.5% |
4.3.1 Personalization of the TTM model
Most of users personalized the TTM models giving audio samples as input. They all started from the off-the-shelf pretrained AudioLDM2 model.
In general, the users really appeciated the possibility of personalizing the model according to their taste. We here report some comments from the participants. If the code of the participant is followed by , the sentence has been directly translated from italian (this is valid for the whole paper). P4: ”It further improves the quality of the modeling the task of generating samples from a specific genre. Even if this means that the quality in very different genres is decreased it’s not a big deal, as we can still fall back to the original model or even fine-tune it again for a different task.” P7: ”It is good because it can fit better your own music tastes”. P18: ”The result is more likeable to the user.”
Particular attention was given to the potential of model personalization in crafting a distinct sound signature for the user. P2: ”It could be useful in order to save your general signature sound and use it to create new sounds starting from them.” P13: ”It can provide signature sounds that only belongs to the artists.”
However, issues concerning the copyright related to the TTM models personalization were also raised. P17 thinks that ”the riff generated after customization is too similar to that of the input data, at risk of plagiarism.”
Users were particulary impressed by the personalization ability of the models with respect to rhythms and percussive sounds. P3*: ”In terms of tone generation of the customization, the drums are better reproduced.” P10*: ”With the basic model, the audio generated does not have a rhythm or cadence, while with the customized model there is more rhythm and the tempo is better marked.” P14* noticed that during the personalization ”the sound is very dissonant (bad even on a harmonic level), but the percussive sounds are well reproduced.” P21*: ”It is very good at personalizing the rhythm.”.
Several users expressed interest in personalizing the same model repeatedly by inserting multiple text-audio concepts into it, to verify whether the model could distinguish multiple concepts simultaneously or if this would cause an exclusive overlap of concepts. P2* states that ”Fine-tuning the model greatly increases the quality in the generation of one specific genre, but it also lowers the average quality when generating music from completely different genres.” P5 claims that ”The fine-tuning overlay only works individually.” P13 expressed the desire to train the model with three different text-audio pairs (bass, drums, and synth lead sound) to generate a track containing all three sound classes. After completing the test, the participant stated that ”the model can reproduce the desired sounds singular, it fails to join all three of them effectively within a composition. I noted that there is always one of the three instruments that prevail over the others.” The same participant also noticed that when attempting to combine multiple personalizations, there was an overlap of sounds and a prevalence of the most recently customized audio dataset.
4.4 Consistency, Expectations, and Quality
Table 2 reports the results related to the Model evaluation survey. Each participant was asked to take the survey after every interaction with the model. The survey aimed to gain insight into participants’ perceptions regarding three aspects: the consistency between the generated audio files and input prompts, the consistency between the generated audio samples and user expectations, and the overall quality of the audio(s) output concerning the user expectations. Only a limited part of the participants considered the overall generation to be entirely aligned with their expectations, both in terms of content and quality. The same holds for the consistency between audio and input prompts. The majority of participants rated all three aspects between and (on a Likert scale from 1 to 5), suggesting that they perceived the generated audio to align reasonably well with their expectations and the input prompts provided. Several participants were positively surprised by the audio generated, as it exceeded their expectations. However, it is important to acknowledge that the subjective nature of the expectation scale presents a limitation in this study. We did not collect pre-generated user expectation data, but we intend to address this in future experimental designs.
Many participants were surprised by some of the generated audio files: some of them expected from the TTM model the ability to sing a textual prompt (P5, P8, P11), a spoken dialogue with understandable language (P11, P24), or a single instrument instead of an entire musical beat (P7, P13). Additionally, they noted that the timbre of guitars and drums closely matched their expectations, resembling the authentic sound of these instruments. Many expected the model to have the ability to adapt an instrument, sound, or melody to a certain different musical style (i.e. performing style transfer), but this was not possible by design with the model/techniques considered. This indicates, in our view, that users should be informed and guided about the generation capabilities of TTM techniques.
4.5 Integration of TTM models in the creative process
At the end of the interaction process, we asked participants to complete a final questionnaire, containing more general questions regarding the experience. Between them, there were some open and multiple-choice questions focusing on the integration of TTM models in the creative process. In Table 3 we report answers to questions evaluated on a Likert scale, while in the remainder of the section, we discuss the open questions contained in the final questionnaire.
In question 8, we asked if they would include this workflow in their music creative process and how. The of participants selected I would only use it for inspiration purpose, while the chose I would take the audio generated and modify it in post-generation. Other participants expressed different comments. P17* asserted: ”At the moment, no, because the program does not process music according to my specifications”, suggesting the need for higher controllability. P10* commented: ”I would use it for specific exercises on rhythm and improvisation based on the generated audio”, which is a type of application that was not considered when we designed the questionnaire.
In the same questionnaire, we asked in what context the users would use the audio generated by the model both in the case of original and personalized models. The majority of the participants provided expected answers related to the use of samples/loops/beats for music production, for sound design, or as an inspiration tool. Others provided original suggestions, probably related to their personal work and creative context, such as the use for foley sounds generation, or P13*: ”during dance lectures”, and P14*: ”I would use it to build musical tracks for dance choreography”. The latter two are particularly interesting, suggesting that when considering the possible applications of TTM models, it is necessary to also consider the background of the specific user, which might desire a final applications not imagined beforehand by the designers of the models. Some participants, referring to the personalized model, suggested that ”The personalized model must adhere more closely to the specifications to be used” (P17*), or ”I would not use it because it did not meet my expectations” (P10*). This suggests that the model’s adherence to the desired objectives is important to users.
P14 suggested: ”I would use it to write music from a particular sub-genre”. This is interesting since most models can generate music belonging to mainstream genres, while some users might need to generate, instead, music belonging to certain sub-genres, which are less represented in commonly available music datasets for training the model.
Important insights considered the role of TTM models in the process of democratization of AI tools for music production. P1: ”I believe that integration of artificial intelligence can give a boost to music production, providing more opportunities for young artists.”. However, the potentials and risks of TTM models are well clear to some participants. P5: ”I think it has the potential to let musicians get realistic sounds with very little effort, but it could also be a bit risky as it could essentially allow anyone to copy any style in very little time, possibly harming the original creator (e.g. by providing much cheaper copies).
5. Discussion
In this section, we summarize the research outcomes derived from the proposed study by offering broad insights into users’ perceptions of TTM models and their role in the creative process.
High-quality generated samples are not all that matters. Sometimes they are expected by users, especially in music production, but there are contexts where they are not essential. In music creation, for instance, decent audio samples from TTM models can often suffice to stimulate creativity and be directly integrated into musicians’ workflows. Even if most of the participants did not rate the audio quality as the best (most of them rated out of the audio quality of the generated sample - Table 2), they stated that they would use the generated sounds directly, as they are, for inspiration, beat-making, music production, and even for crafting tracks for dance practice. This suggests that users which use TTM models to generate sounds for inspirational purposes may not need high-quality sounding generated audio. Instead, they might prioritize the creativity and novelty of the generated sounds over quality.
Context and subjective perception are important when evaluating TTM models. Analyzing the perception of TTM systems based on different contexts and with participants with different backgrounds made clear how challenging is to objectively evaluate TTM models. This finding is consistent with results presented in (Vinay and Lerch,, 2022) regarding the limitations of objective metrics in evaluating generative audio systems. Metrics commonly used to evaluate generative models, such as the Fréchet Audio Distance (Kilgour et al.,, 2019), while informative, often fail to capture the nuances of user interactions and preferences. Assessing usability requires a thorough understanding of how these systems are used and perceived by end-users, as also highlighted in (Chu et al.,, 2022) in the context of automatic symbolic music generation models. Determining the usefulness of TTM model outputs is not straightforward when relying only on objective metrics, suggesting the need for novel and complementary subjective evaluative approaches that also consider usability, user-centered experience, and contextual relevance.
TTM models contribute to democratizing the music creation process, but raise concerns regarding copyright protection. Some participants emphasized how the model and the personalization technique explored in this study empower individuals to create their unique sound signatures, enabling a user-centered audio generation. However, concerns arise about control and copyright protection. As also pointed out by some participants, with personalization techniques becomes easy to incorporate others’ signature sounds into the generative model, making them vulnerable to unauthorized use. While music AI-based systems democratize music creation, making it accessible to a wider user base, there is a crucial need to address ethical guidelines and copyright laws regarding training datasets and the use of generated sounds (Ren et al.,, 2024; Franceschelli and Musolesi,, 2022).
Controllability and editability are at the forefront of requested features. Some limitation of the model considered in this study concerns the lack of editability, allowing the users to actively shape the direction of the audio generation process based on their choices. Participants expressed the desire to have more control on the generation process, for example, to modify only a specific section of the generated music track. It is important then, that the next generation of TTM models will also include proper user interfaces, which allow the users to set and modify specific parameters to guide the generation of the desired audio. Moreover, the graphical interfaces cannot be the same for every user but need to be contextualized according to the final users and their final tasks. In fact, some of these features have been integrated in the just released models (at the time of writing), such as Stable Audio 2.0 (from Stability AI) and Project Music GenAI Control (from Adobe).
6. Conclusions and future work
This study proposes to investigate the interaction between users and Text-to-Music generative models, to understand how users would employ and perceive these tools. We developed a simple interface to engage with users. Through the interface, participants of the study could generate audio from textual prompts, using a selected TTM model, and/or personalize the proposed model to generate specific sounds based on their preferences. During the study, through the use of questionnaires, we asked them specific questions related to the interaction with the interface and to TTM models at large. Results show that these models have several application opportunities, not only in the music creative process. Particular attention was given to the opportunities of the model’s personalization. At the same time, worries about plagiarism and unauthorized use of generated personalized sounds were raised. Moreover, despite the assertion that we have yet to develop a fully functional AI music tool capable of satisfying both technical and contextual considerations, as well as its final application, the majority of participants strongly asserted that these tools have significant growth potential and offer a wide range of applications across various domains. Future works aim to include these results in TTM generative models and develop a better interface to allow further interaction and control for the final users.
7. Ethics Statement
We acknowledge that diversity of nationality and education level of the pool of participants is limited. We however believe that their interest and knowledge and AI makes the results and their comments and suggestion interesting to the wider research community. Participants were able to use samples of their own to personalize the model. These samples might also contain music under copyright, we would like to stress that no music generated through the model was made available publicly.